OpenStudy (anonymous):

What information is given when f(x) is used instead of y in the representation of a mathematical equation, such as f(x) = 2x + 3? How is this information useful?

OpenStudy (anonymous):

Its useful one to say that an equation is a function

OpenStudy (anonymous):

For ex \[\Large y=\pm \sqrt{1-x^2}\] is written for y= something right? But it's a not a function, it just shows how y and x are related

OpenStudy (anonymous):

This notation is also useful becuse you can have something like f(5)=2(5)+3, it shows that at 5 what the left side will be.